OpenStudy (anonymous):

Can someone help me set up this diff eq? I will write it in math below. I tried to use t=uy and dt=udy+y but really didn't get anything workable.

OpenStudy (anonymous):

\[3(1+t^2)\frac{ dy }{ dt }=2ty(y^3-1)\]

OpenStudy (phi):

I would re-arrange it as \[ \frac{3}{y(y^3-1)} dy=\frac{2t}{t^2+1} dt \] the right hand side is of the form du/u with u= (t^2+1) the left side can be separated using partial fractions (see http://www.purplemath.com/modules/partfrac.htm)